what is the slope of 6y=4x-6

Found 2 solutions by mananth, rosey51235:
Answer by mananth(16949)

6y=4x-6
y=4/6 x -1
slope = 4/6 = 2/3
2/3

Answer by rosey51235(30)

Hello!
First, you need to know that finding the slope of a linear equation is very easy, so long as it's in the right form. This is already in the proper form, which is y=mx+b (or in this case, + a - number, so just -b).
So, since it's already in the correct form, you just have to solve it by getting y alone.
Start by dividing everything by 6, which turns 6y=4x-6 into y=2/3x-1...
In this case, leave the 2/3 as a fraction, since slope is rise over run...so in this case the slope is rise 2, run 3. 2/3 (because in the equation y=mx+b m= the slope, and b= the y intercept, where the line formed by the equation should cross the y-axis.)
Hope I helped! :)
Rosey